# Harnessing the edge of chaos for combinatorial optimization

### v2

@@ Abstract
Nonlinear dynamical systems with continuous variables can be used for solving combinatorial optimization problems with discrete variables. In particular, numerical simulations of them can be used as heuristic algorithms with a desirable property, namely, parallelizability, which allows us to execute them in a massively parallel manner using cutting-edge many-core processors, leading to ultrafast performance. However, the dynamical-system approaches with continuous variables are usually less accurate than conventional approaches with discrete variables such as simulated annealing. To improve the solution accuracy of a representative dynamical system-based algorithm called simulated bifurcation (SB), which was found from classical simulation of a quantum nonlinear oscillator network exhibiting quantum bifurcation, here we generalize it by introducing nonlinear control of individual bifurcation parameters and show that the generalized SB (GSB) can achieve almost 100% success probabilities for some large-scale problems. As a result, the time to solution for a 2,000-variable problem is shortened to 10 ms by a GSB-based machine, which is two orders of magnitude shorter than the best known value, 1.3 s, previously obtained by an SB-based machine. To examine the reason for the ultrahigh performance, we investigated chaos in the GSB changing the nonlinear-control strength and found that the dramatic increase of success probabilities happens near the edge of chaos. That is, the GSB can find a solution with high probability by harnessing the edge of chaos. This finding suggests that dynamical-system approaches to combinatorial optimization will be enhanced by harnessing the edge of chaos, opening a broad possibility to tackle intractable combinatorial optimization problems by nature-inspired approaches.
@@ III Performance of the GbSB
To evaluate the performance of the GbSB, we first solved a standard benchmark problem named K2000, a 2,000-node MAX-CUT problem [51] equivalent to a 2,000-spin Ising problem with all-to-all connectivity ($J_{i,j}\in\{\pm 1\}$[8, 20, 22, 24, 27, 35]. The best known cut value of K2000 is 33,337 reported in ref. 24. Figure 1a shows the success probability, $P_{\mathrm{S}}$, of finding the best known value by the GbSB with various values of $M$ and $A$. (See Appendix A for the detailed settings of the other parameters.) It thus turns out that the GbSB with appropriate values of $M$ and $A$ can find the best known value of K2000 with almost 100% probability. To the best of our knowledge, this is the highest success probability of this problem.
To verify that the ultrahigh performance of the GbSB holds for problems other than K2000, we also solved randomly generated 100 instances of the 700-spin Ising problem with all-to-all connectivity ($J_{i,j}\in\{\pm 1\}$}[24, 30] and the first ten instances of the well-known MAX-CUT benchmark set called G-set [8, 18, 19, 20, 24, 29, 31, 33, 34, 35, 48], which are equivalent to the 800-spin Ising problem with sparse connectivity ($J_{i,j}\in\{0,-1\}$} or $\{0,\pm 1\}$, see Fig. 2). Similarly to the case of K2000, we obtained success probabilities exceeding 90% for 32 instances of the 700-spin problem and two G-set instances (G3 and G6). (See Tables 1 and 2 for the detailed results and settings.) Figures 1b and 1c show the corresponding results of two of them to Fig. 1a. Thus, the ultrahigh performance of the GbSB holds not only for K2000 but also for other large-scale problems.
@@ IV Edge-of-chaos effect in the GbSB
The lower figures in Figs. 1d–1f show the results of $\delta(t_{M})$ ($\delta$ at the final time) corresponding to Figs. 1a–1c, respectively. In all the three cases, $\delta(t_{M})\simeq 0$} when A=0{A=0}, indicating regular dynamics, and $\delta(t_{M})$ converges to $1/\sqrt{2}$} (dotted lines) as $A$ increases, indicating the chaos due to the nonlinear control. The regions highlighted in blue in Figs. 1d–1f show the condition that the success probabilities become particularly high, as shown in the upper figures. We thus find that the success probabilities become high near the edge of chaos (between regular and chaotic regions). This result suggests that weakly chaotic processes occurring near regular dynamics may assist avoiding trapping at local minima and finding optimal solutions in the GbSB. This chaos-assisted solution search may be essentially different from conventional approaches with random noises such as SA, because the dramatic increase of the success probabilities, in particular, almost 100% success probabilities for the large-scale problems, has not been achieved by SA.
@@ V No influence of discretization
We also investigated the dependence of the performance on the time step $\Delta_{\mathrm{t}}$. The results in Figs. 1g–1i show that the notable behavior of the GbSB is independent of $\Delta_{\mathrm{t}}$. This result suggests that the high performance of the GbSB harnessing the edge of chaos comes not from artifacts due to the discretization for numerical simulations, but from the original Hamiltonian equations of motion with the walls and the nonlinear control of individual bifurcation parameters.
@@ VI GbSB-based machine with an FPGA
To demonstrate high parallelizability of the GbSB, we designed a highly parallel custom circuit for the GbSB to minimize the execution time and then instantiated a 2,048-spin GbSB machine (GbSBM) with an FPGA (Fig. 2). Each individual bifurcation parameter $p_{i}(t)$ newly introduced in the GbSB is updated depending only on its corresponding position $x_{i}(t)$ without depending on the other oscillators. Hence, the GbSB is highly parallelizable, as conventional SBs are [22, 23, 24, 25, 26]. In the computation for a GbSB time-evolution step, the many-body interaction computation, $\sum_{j=1}^{N}J_{i,j}x_{j}(t_{m})$, in Eq. (3) is the most computationally intensive with a computational complexity of $\Theta(N^{2})$, while the remaining parts including updating $p_{i}(t)$, $x_{i}(t)$, and $y_{i}(t)$, correctively referred to as PXY, have a computational complexity of $\Theta(N)$. We accelerate the many-body interaction computation by spatial parallelization (referred to as JX in Fig. 2a), which is characterized by three parallelization parameters of $P_{\mathrm{r}}$, $P_{\mathrm{c}}$, and $P_{\mathrm{b}}$ (Figs. 2a and 2b). The PXY part is accelerated by temporal parallelization (Fig. 2a). The number of the clock cycles per GbSB time-evolution step is expressed as
@@
where $\lambda$ is the maximum circuit latency for circulative paths (See Appendix B for details). The circuit architecture accelerates the many-body interaction computation by $P_{\mathrm{r}}P_{\mathrm{c}}P_{\mathrm{b}}$ and allows that $P_{\mathrm{r}}P_{\mathrm{c}}P_{\mathrm{b}}$ can be larger than the number of oscillators $N$.
@@ VII Summary and outlook
We have generalized bSB by introducing nonlinear control of individual bifurcation parameters such that the oscillators in the bSB do not stick to walls. We have found that the success probabilities of the generalized bSB (GbSB) for large-scale problems dramatically increase and sometimes approach almost 100%. Developing a GbSB-based machine with an FPGA, we have demonstrated one or two orders of magnitude improvements of times to solution for the large-scale problems compared with a previously developed SB-based FPGA machine. By investigating chaos in the GbSB, we have also found that the surprising increase of the success probabilities occurs near the edge of chaos. This result suggests that the GbSB can achieve the ultrahigh performance by harnessing the edge of chaos.
Our results open broad and interesting possibilities as follows. In this work, we have focused on bSB and found the edge-of-chaos effect. It is thus interesting whether or not other dynamical-system approaches, such as aSB [22], dSB [24], neural networks [4, 5], coherent Ising machines [6, 7, 8, 9, 10, 11], and oscillator-based Ising machines [33, 34], exhibit similar effects by similar nonlinear control. The GbSB has an additional parameter, namely, the nonlinear-control strength, which requires additional tuning for achieving high performance. The edge of chaos will be helpful to find the best condition for the nonlinear-control strength. The GbSB could not improve G9, one of the G-set instances, compared with dSB. This suggests that we may need a hybrid algorithm including the GbSB and others for solving various problems. In general, the SBs can achieve high performance for dense-connectivity problems [3, 69]. On the other hand, a dynamical-system approach that can achieve ultrahigh performance for sparse-connectivity problems has recently been reported [70]. The combination of them will be promising to tackle both kinds of problems.
@@ Appendix A Parameter settings in the GbSB
As for initial conditions, $x_{i}(0)$ was set to a random value between 1-1 and 1, except the simulations to evaluate $\delta(t_{M})$ in Figs. 1d–1f, and $y_{i}(0)$ was set to zero.
Appendix B FPGA implementation of GbSB
Figure 4a shows the timing chart of the operation of JX and PXY modules. As shown in Fig. 4b, a JX module processes $P_{\mathrm{r}}P_{\mathrm{c}}$ multiply-accumulation operations per clock cycle and then produce $P_{\mathrm{r}}$ numbers of $\Delta y_{i}$ after $N/P_{\mathrm{c}}$ cycles (hereafter, a phase), and there are $P_{\mathrm{b}}$ JX modules. A PXY module shown in Fig. 2a takes as input a $\Delta y_{i}$ datum per clock cycle and processes $P_{\mathrm{b}}P_{\mathrm{r}}/P_{\mathrm{c}}$ numbers of $\Delta y_{i}$ data per phase, and there are $P_{\mathrm{c}}$ PXY modules. The operation of a PXY module is overlapped with that of JX modules in time domain (Fig. 4a). The maximum circuit latency for circulative paths is denoted by $\lambda$, meaning that the system must wait for $\lambda$ to guarantee that the final datum has been stored in memory after the final iteration begins. Thus, the number of the clock cycles per GbSB time-evolution step, $N_{\mathrm{cyc}}$, is expressed by Eq. (9).

### v4

# Edge-of-chaos enhanced quantum-inspired algorithm for combinatorial optimization

@@ Abstract
Nonlinear dynamical systems with continuous variables can be used for solving combinatorial optimization problems with discrete variables. Numerical simulations of them are also useful as heuristic algorithms with a desirable property, namely, parallelizability, which allows us to execute them in a massively parallel manner, leading to ultrafast performance. However, the dynamical-system approaches with continuous variables are usually less accurate than conventional approaches with discrete variables such as simulated annealing. To improve the solution accuracy of a quantum-inspired algorithm called simulated bifurcation (SB), which was found from classical simulation of a quantum nonlinear oscillator network exhibiting quantum bifurcation, here we generalize it by introducing nonlinear control of individual bifurcation parameters and show that the generalized SB (GSB) can achieve surprisingly high performance, namely, almost 100% success probabilities for some large-scale problems. As a result, the time to solution for a 2,000-variable problem is shortened to 10 ms by a GSB-based machine, which is two orders of magnitude shorter than the best known value, 1.3 s, previously obtained by an SB-based machine. To examine the reason for the ultrahigh performance, we investigated chaos in the GSB changing the nonlinear-control strength and found that the dramatic increase of success probabilities happens near the edge of chaos. That is, the GSB can find a solution with high probability by harnessing the edge of chaos. This finding suggests that dynamical-system approaches to combinatorial optimization will be enhanced by harnessing the edge of chaos, opening a broad possibility to tackle intractable combinatorial optimization problems by physics-inspired approaches.
@@ II Generalized ballistic simulated bifurcation (GbSB)
Figure 1: Numerical results of the GbSB. (a) Success probability, $P_{\mathrm{S}}$, of finding the best known cut value of K2000 by the GbSB with $\Delta_{t}=1.25$} and various values of $M$ and $A$. See Appendix A for the setting of $c$. The lower figure is the enlargement of the part enclosed by the dashed rectangle in the upper figure. $P_{\mathrm{S}}$ was evaluated by 1,000 times repetition with different initial conditions. (b) and (c) Results corresponding to (a) for an instance of the 700-spin Ising problem and G6, respectively. The time step is set as $\Delta_{t}=1.25$} and $D_{t}=1.25$} [see Eq. (13) in Appendix A], respectively. See Appendix A for the setting of $c$. (d) $P_{\mathrm{S}}$ in a (upper figure) and normalized distance, $\delta(t_{M})$, at the final time (lower figure) for K2000 when M=21,500{M=21,500} in (a). $\delta(t_{M})$ was evaluated by averaging 100 results with different initial conditions. The regions highlighted in blue show the condition that $P_{\mathrm{S}}$ becomes particularly high. The dotted line in the lower figure shows $\delta(t_{M})=1/\sqrt{2}$, which indicates chaos. (e) and (f) Results corresponding to (d) for the 700-spin instance when M=3,000{M=3,000} in (b) and G6 when M=1,000{M=1,000} in (c), respectively. (g–i) $P_{\mathrm{S}}$ for K2000, the 700-spin instance, and G6, respectively, for different values of the time step $\Delta_{\mathrm{t}}$ and different final times $t_{M}=\Delta_{\mathrm{t}}M$. AA was set to 0.2, 0.25, and 0.2, respectively. (j–l) $\delta(t_{M})$ corresponding to (g–i), respectively.
@@ III Performance of the GbSB
To evaluate the performance of the GbSB, we first solved a standard benchmark problem named K2000, a 2,000-node MAX-CUT problem [51] equivalent to a 2,000-spin Ising problem with all-to-all connectivity ($J_{i,j}\in\{\pm 1\}$[8, 20, 22, 24, 27, 35]. In this work, we evaluate the cut value only at the final time, because its calculation is time-consuming. (We confirmed that the present results remain unchanged even if we evaluate the cut value at all the time steps and select the best one.) The best known cut value of K2000 is 33,337 reported in ref. 24. Figure 1(a) shows the success probability, $P_{\mathrm{S}}$, of finding the best known value by the GbSB with various values of $M$ and $A$. [See Appendix A for the detailed settings of the other parameters. Also see Appendix B for average cut values corresponding to Figs. 1(a–c).] It thus turns out that the GbSB with appropriate values of $M$ and $A$ can find the best known value of K2000 with almost 100% probability. To the best of our knowledge, this is the highest success probability of this problem. (In Fig. 1(b), we can see a vertical stripe around M=2700{M=2700} where $P_{\mathrm{S}}$ becomes relatively low. This is not an artifact but its underlying cause is difficult to explain. This may be a unique circumstance of this instance.)
To verify that the ultrahigh performance of the GbSB holds for problems other than K2000, we also solved randomly generated 100 instances of the 700-spin Ising problem with all-to-all connectivity ($J_{i,j}\in\{\pm 1\}$}[24, 30] and the first ten instances of the well-known MAX-CUT benchmark set called G-set [8, 18, 19, 20, 24, 29, 31, 33, 34, 35, 48], which are equivalent to the 800-spin Ising problem with sparse connectivity ($J_{i,j}\in\{0,-1\}$} or $\{0,\pm 1\}$, see Fig. 2). Similarly to the case of K2000, we obtained success probabilities exceeding 90% for 32 instances of the 700-spin problem and two G-set instances (G3 and G6). (See Tables 1 and 2 for the detailed results and settings.) Figures 1(b)(c) show the corresponding results of two of them to Fig. 1(a). Thus, the ultrahigh performance of the GbSB holds not only for K2000 but also for other large-scale problems.
@@ IV Edge-of-chaos effect in the GbSB
The lower figures in Figs. 1(d–f) show the results of $\delta(t_{M})$ ($\delta$ at the final time) corresponding to Figs. 1(a–c), respectively. In all the three cases, $\delta(t_{M})\simeq 0$} when A=0{A=0}, indicating regular dynamics, and $\delta(t_{M})$ converges to $1/\sqrt{2}$} (dotted lines) as $A$ increases, indicating the chaos due to the nonlinear control. The regions highlighted in blue in Figs. 1(d–f) show the condition that the success probabilities become particularly high, as shown in the upper figures. We thus find that the success probabilities become high near the edge of chaos (between regular and chaotic regions). This result suggests that weakly chaotic processes occurring near regular dynamics may assist avoiding trapping at local minima and finding optimal solutions in the GbSB. This chaos-assisted solution search may be essentially different from conventional approaches with random noises such as SA, because the dramatic increase of the success probabilities, in particular, almost 100% success probabilities for the large-scale problems, has not been achieved by SA.
@@ V No influence of discretization
We also investigated the dependence of the performance on the time step $\Delta_{\mathrm{t}}$. The results in Figs. 1(g–i) show that the notable behavior of the GbSB is independent of $\Delta_{\mathrm{t}}$. This result suggests that the high performance of the GbSB harnessing the edge of chaos comes not from artifacts due to the discretization for numerical simulations, but from the original Hamiltonian equations of motion with the walls and the nonlinear control of individual bifurcation parameters.
Figures 1(g–i) also show that $P_{S}$ does not monotonically increase with respect to the number of steps $M$, unlike SA. To investigate the reason for this, we checked the normalized distance at the final time $\delta(t_{M})$ corresponding to Figs. 1(g–i), the results of which are shown in Figs. 1(j–l). Interestingly, Fig. 1(j) suggests that $P_{S}$ increases around the “valley of chaos.” This is also the case in Figs. 1(k) and 1(l), though they are less clear. This is natural in the sense that the butterfly effect becomes more evident for longer time evolution. Thus, a too large $M$ (more precisely $t_{M}$) may result in chaos and consequently $P_{S}$ may become lower.
@@ VI GbSB-based machine with an FPGA
Figure 2: GbSB-based machine with an FPGA. (a) Block diagram of the GbSB-based machine, which has a circulative structure corresponding to iteration of updating position and momentum: $y_{i}(t_{m+1})=y_{i}(t_{m})+\delta y_{i}(t_{m})+\Delta y_{i}(t_{m})$ and $x_{i}(t_{m+1})=x_{i}(t_{m})+\delta x_{i}(t_{m})$, where $\Delta y_{i}(t_{m})=\mathrm{JX}[\mathbf{J}_{i},\mathbf{x}(t_{m})]=c\Delta_{\mathrm{t}}\sum_{j=1}^{N}J_{i,j}x_{j}(t_{m})$, $p_{i}(t_{m+1})=\mathrm{FP}[p_{i}(t_{m}),x_{i}(t_{m})]=p_{i}(t_{m})-[1-Ax_{i}(t_{m})^{2}]p_{i}(t_{m})/(M-m)$, $\delta y_{i}(t_{m})=\mathrm{FX}[p_{i}(t_{m+1}),x_{i}(t_{m})]=-p_{i}(t_{m+1})x_{i}(t_{m})\Delta_{\mathrm{t}}$, and $\delta x_{i}(t_{m})=\mathrm{FY}[y_{i}(t_{m+1})]=y_{i}(t_{m+1})\Delta_{\mathrm{t}}$. The JX, FP, FX, FY modules in the block diagram correspond to the JX, FP, FX, FY functions, and the wall module executes the following conditional operation: $(x_{i}(t_{m+1}),y_{i}(t_{m+1}))\leftarrow\left\{\begin{array}[]{ll}(\mathrm{sgn}(x_{i}(t_{m+1})),0),&\mathrm{if}~|x_{i}(t_{m+1})|>1\\ (x_{i}(t_{m+1}),y_{i}(t_{m+1})),&\mathrm{otherwise}\end{array}\right$.. The computational precision of the JX module is 16-bit fixed point, while that for the remaining modules is 32-bit floating point. (b) Block diagram of a many-body interaction (JX) module. The parallelization parameters ($P_{\mathrm{r}}$, $P_{\mathrm{c}}$, and $P_{\mathrm{b}}$) are illustrated in a and b. (c) Layout of the circuit modules in the FPGA, where the routing congestion is shown as a heatmap with the regions for the JX, PXY, and BSP (board support package) modules being indicated by dashed lines.
Figure 3: Times to solution (TTSs) with a GbSB-based FPGA machine. The second row (N=700{N=700}) is the median of the TTSs for 100 random instances of the 700-spin Ising problem. See Appendix A for the detailed parameter settings and the GbSB-based machine (GbSBM). The results with the dSB-based machine (dSBM) are cited from ref. 24.
To demonstrate high parallelizability of the GbSB, we designed a highly parallel custom circuit for the GbSB to minimize the execution time and then instantiated a 2,048-spin GbSB machine (GbSBM) with an FPGA (Fig. 2). Each individual bifurcation parameter $p_{i}(t)$ newly introduced in the GbSB is updated depending only on its corresponding position $x_{i}(t)$ without depending on the other oscillators. Hence, the GbSB is highly parallelizable, as conventional SBs are [22, 23, 24, 25, 26]. The increase in computational complexity resulting from introducing individual parameters $p_{i}(t)$} [Eq. (7)] instead of a common parameter p(t){p(t)} [Eq. (6)] is $\Theta(N)$}. In the computation for a GbSB time-evolution step, the many-body interaction computation, $\sum_{j=1}^{N}J_{i,j}x_{j}(t_{m})$, in Eq. (3) is the most computationally intensive with a computational complexity of $\Theta(N^{2})$, while the remaining parts including updating $p_{i}(t)$, $x_{i}(t)$, and $y_{i}(t)$, correctively referred to as PXY, have a computational complexity of $\Theta(N)$. We accelerate the many-body interaction computation by spatial parallelization [referred to as JX in Fig. 2(a)], which is characterized by three parallelization parameters of $P_{\mathrm{r}}$, $P_{\mathrm{c}}$, and $P_{\mathrm{b}}$ [Figs. 2(a)(b)]. The PXY part is accelerated by temporal parallelization [Fig. 2(a)]. The number of the clock cycles per GbSB time-evolution step is expressed as
@@
where $\lambda$ is the maximum circuit latency for circulative paths (See Appendix C for details). The circuit architecture accelerates the many-body interaction computation by $P_{\mathrm{r}}P_{\mathrm{c}}P_{\mathrm{b}}$ and allows that $P_{\mathrm{r}}P_{\mathrm{c}}P_{\mathrm{b}}$ can be larger than the number of oscillators $N$.
@@ VII Summary and outlook
We have generalized bSB by introducing nonlinear control of individual bifurcation parameters such that the oscillators in the bSB do not stick to walls. We have found that the success probabilities of the generalized bSB (GbSB) for large-scale problems dramatically increase and sometimes approach almost 100%. Developing a GbSB-based machine with an FPGA, we have demonstrated one or two orders of magnitude improvements of times to solution for the large-scale problems compared with a previously developed SB-based FPGA machine. By investigating chaos in the GbSB, we have also found that the surprising increase of the success probabilities occurs near the edge of chaos. This result suggests that the GbSB can achieve the ultrahigh performance by harnessing the edge of chaos. The present study on chaos is based on numerical experiments. Theoretical studies on the present system are desirable but left for future work. Also, the above conclusion is based on numerical study on 111 specific instances, and therefore investigating more instances from a wider range of problem classes is also desirable.
Our results open broad and interesting possibilities as follows. In this work, we have focused on bSB and found the edge-of-chaos effect. It is thus interesting whether or not other dynamical-system approaches, such as aSB [22], dSB [24], neural networks [4, 5], coherent Ising machines [6, 7, 8, 9, 10, 11], and oscillator-based Ising machines [33, 34], exhibit similar effects by similar nonlinear control. The GbSB has an additional parameter, namely, the nonlinear-control strength, which requires additional tuning for achieving high performance, together with tuning of the number of time steps. Although we have not included the time to optimize these parameters in computation times in this work, as in previous studies, the optimization time is not negligible for practical applications. To optimize the parameters quickly, our finding that good parameter values may be found around the edge of chaos will be helpful. The GbSB could not improve G9, one of the G-set instances, compared with dSB. This result suggests the limitation of the power of GbSB. While it is important to clarify the range of applicability of GbSB, this result suggests that we may need a hybrid algorithm combining GbSB with others for solving a wide range of problems. In general, the SBs can achieve high performance for dense-connectivity problems [3, 69]. On the other hand, a dynamical-system approach that can achieve ultrahigh performance for sparse-connectivity problems has recently been reported [70]. The combination of them will be promising to tackle both kinds of problems.
@@ Appendix A Parameter settings in the GbSB
As for initial conditions, $x_{i}(0)$ was set to a random value between 1-1 and 1, except the simulations to evaluate $\delta(t_{M})$ in Figs. 1(d–f), and $y_{i}(0)$ was set to zero.
Appendix B Average cut values in Figs. 1(a–c)
Figures 1(a–c) show the success probability $P_{S}$ to obtain the best known value, where $P_{S}\simeq 0$ in a wide range of the parameters. Here we show the average cut value $C_{\mathrm{ave}}$ normalized by the best know value $C_{\mathrm{best}}$ in Fig. 4. We can confirm the gradual improvement of $C_{\mathrm{ave}}$ by GbSB.
Figure 4: Average cut value $C_{\mathrm{ave}}$ normalized by the best know value $C_{\mathrm{best}}$ corresponding to Figs. 1(a–c), respectively.
Appendix C FPGA implementation of GbSB
Figure 5(a) shows the timing chart of the operation of JX and PXY modules. As shown in Fig. 5(b), a JX module processes $P_{\mathrm{r}}P_{\mathrm{c}}$ multiply-accumulation operations per clock cycle and then produce $P_{\mathrm{r}}$ numbers of $\Delta y_{i}$ after $N/P_{\mathrm{c}}$ cycles (hereafter, a phase), and there are $P_{\mathrm{b}}$ JX modules. A PXY module shown in Fig. 2a takes as input a $\Delta y_{i}$ datum per clock cycle and processes $P_{\mathrm{b}}P_{\mathrm{r}}/P_{\mathrm{c}}$ numbers of $\Delta y_{i}$ data per phase, and there are $P_{\mathrm{c}}$ PXY modules. The operation of a PXY module is overlapped with that of JX modules in time domain [Fig. 5(a)]. The maximum circuit latency for circulative paths is denoted by $\lambda$, meaning that the system must wait for $\lambda$ to guarantee that the final datum has been stored in memory after the final iteration begins. Thus, the number of the clock cycles per GbSB time-evolution step, $N_{\mathrm{cyc}}$, is expressed by Eq. (9).